OpenStudy (anonymous):

find an equation of the tangent line at point given: f(x)=8/x^2+4; (2,1)

myininaya (myininaya):

Have you differentiated f yet?

OpenStudy (anonymous):

not sure how

OpenStudy (anonymous):

would i have to plug in the point into the point slope form ? or

myininaya (myininaya):

\[f(x)=\frac{8}{x^2+4} \text{ or } f(x)=\frac{8}{x^2}+4 ?\]

myininaya (myininaya):

I think you mean the first one because (2,1) isn't on the second graph

myininaya (myininaya):

To differentiate f you need to know the quotient rule

myininaya (myininaya):

Or at least the product rule and chain rule

OpenStudy (anonymous):

yeah the first one

myininaya (myininaya):

Do you know the quotient rule?

OpenStudy (anonymous):

f'g-fg'/g^2 ?

myininaya (myininaya):

\[f'(x)=\frac{(8)'(x^2+4)-8(x^2+4)'}{(x^2+4)^2}\]

myininaya (myininaya):

so you need to know the derivative of 8 and the derivative of x^2+4 to finish that

OpenStudy (anonymous):

okie dokie !

myininaya (myininaya):

once you are done with that replace the x's with 2' to find f'(2) this will be the slope of the tangent line of f at (2,1)

myininaya (myininaya):

then once you know the slope that is it plug in the point-slope form of a line y-y1=m(x-x1)

myininaya (myininaya):

\[y-1=m(x-2)\]

myininaya (myininaya):

your only job is to find m which is f'(2)